OpenStudy (anonymous):

if the polynomial f(x)=ax^{3}+(a+b)x^{2}+(a+2b)x+1is divisible by x+1, find the quotient, leaving its coefficients in term of a only. Prove that in this case, the equation f(x) has only one real root if a^{2}-6a+1<0.Hence, show that 3-2sqrt{2}

OpenStudy (anonymous):

\[f(x)=ax^{3}+(a+b)x^{2}+(a+2b)x+1\]divisible by \[x+1\]means\[f(-1)=0\]\[f(-1)=a(-1)^{3}+(a+b)(-1)^{2}+(a+2b)(-1)+1=0\]\[b=1-a\]therefore\[f(x)=ax^{3}+(a+1-a)x^{2}+(a+2(1-a))x+1\]\[f(x)=ax^{3}+x^{2}+(2-a)x+1\]let\[f(x)=(x+1)(ax^{2}+Ax+1)\]compare coefficient of x,\[2-a=1+A\]\[A=1-a\]therefore,\[f(x)=(x+1)(ax^{2}+(1-a)x+1)\]when f(x)=0,\[0=(x+1)(ax^{2}+(1-a)x+1)\]then i stuck here....

myininaya (myininaya):

neoh i'm so becoming your fan no one ever shows their attempt

OpenStudy (anonymous):

thx...

myininaya (myininaya):

i got b=-1-a

myininaya (myininaya):

did i make a mistake there or did you?

OpenStudy (anonymous):

oh no... i make a mistake... but if using the same method i also stuck at same place...

OpenStudy (dumbcow):

its b =1-a its a "+1" at the end

myininaya (myininaya):

you are good neoh i made the mistake

OpenStudy (dumbcow):

looks good last step, set equal to zero for x=-1 \[ax^{2} +(1-a)x+1 = 0\] \[a(-1)^{2} +(1-a)(-1)+1 = 0\] solve for a

OpenStudy (dumbcow):

wait nevermind, scratch that

OpenStudy (dumbcow):

oh you need to show there is only 1 real root. use discriminant < 0

OpenStudy (anonymous):

i already factorize f(x) by x+1, so the quotient can substitute the same zero again or not??

myininaya (myininaya):

\[(1-a)^2-4a<0\]

OpenStudy (anonymous):

\[(1−a)^{2}-4(a)(1)<0\]\[1−2a+a^{2}-4a<0\]\[a^{2}-6a+1<0\]

OpenStudy (dumbcow):

discriminate : b^2 - 4ac \[\rightarrow (1-a)^{2} -4a < 0\] \[a^{2}-6a+1 < 0\]

myininaya (myininaya):

lol yeah!

OpenStudy (anonymous):

dumbcow is right, just show the discriminant being less than 0. discriminant of \[ax^2+(1-a)x+1 \]

OpenStudy (anonymous):

myininaya - I am hurt.. I always do :P

myininaya (myininaya):

what?

OpenStudy (dumbcow):

haha good job neoh...now just use quadratic formula

OpenStudy (anonymous):

I always reply with the procedure.. not just the answer :D Jus Kidding

OpenStudy (anonymous):

\[(-b \pm \sqrt{b^{2}-4ac})\div 2a\]

myininaya (myininaya):

neat problem

myininaya (myininaya):

kind of long but neat

myininaya (myininaya):

i'm going to sleep later guys

OpenStudy (anonymous):

night

OpenStudy (anonymous):

\[(6 \pm \sqrt{(-6)^{2}-4(1)(1)})\div2(1)\]\[3 \pm 2\sqrt{2}\]

OpenStudy (anonymous):

have two roots... the question say 1 root.... how to prove that??

OpenStudy (anonymous):

to complete the problem you need to check the intervals: \[(-\infty, r_1], (r_1,r_2), [r_2, \infty)\] So that if: \[r_1<a<r_2\] there is only one real root to your equation.

OpenStudy (anonymous):

sry, where: \[r_1 = 3-2\sqrt{2}, r_2 = 3+2\sqrt{2}\]

OpenStudy (anonymous):

and try to keep the perspective straight, these roots we found arent for the polynomial x, they are roots of a.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

If there is only 1 real root then the root must be -1. That means that : \[ax^2+(1−a)x+1\] has no real solution. Using the quadratic formula: \[x = \frac{-(1-a) \pm \sqrt{(1-a)^2-4a}}{2a}\] must not be real thus, the thing under the radical must be negative. Thus \[4a > (1-a)^2 = 1-2a+a^2\] \[0>1-6a+a^2\] as required

OpenStudy (anonymous):

ok...